OpenStudy (anonymous):

The graph of y=10.2 is decreasing. True or False?

OpenStudy (loser66):



OpenStudy (anonymous):

Its a stright line parllel to x-axis.....@bless

OpenStudy (anonymous):

Opss sorry.. My bad. 10.2^x

OpenStudy (anonymous):

then false...

OpenStudy (anonymous):

Why? can you please explain. There's a big test tomorrow. I want to understand

OpenStudy (anonymous):

see this link it explains clearly..... http://www.wolframalpha.com/input/?i=y%3D10.2%5Ex

OpenStudy (loser66):

there is a trick, do you want to know?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes @Loser66

OpenStudy (loser66):

to this kind of problem, you just plug x =1, x =2 into the expression, x from 1 to 2 , it means x increases, right? then y = 10.2 for x =1 y = 104.04 for x =2 you see, x \(\uparrow\) y\(\uparrow\) therefore, it increases, while the statement said it decreases so, the statement is FALSE

OpenStudy (anonymous):

what if the x is equals to -1? Then the graph will look like decreasing?

OpenStudy (loser66):

ok, x =-1 y = 1/10.2 x = -2 , y = 1/104.04 BUT from -2 to -1 , x is increasing (-2< -1) and y from 1/104.04 to 1/10.2 is increasing , too ( 1/104.04<1/10.2, right?) just to confirm, whenever it increases, it increases on the whole interval,

OpenStudy (anonymous):

@Loser66 thank u so much !!

OpenStudy (loser66):

got it?? hehehe, glad to hear

OpenStudy (anonymous):

Yes :)